# Med-R2: Crafting Trustworthy LLM Physicians via Retrieval and Reasoning of Evidence-Based Medicine

### v4

@@ Abstract
Large Language Models (LLMs) have exhibited remarkable capabilities in clinical scenarios. Despite their potential, existing works face challenges when applying LLMs to medical settings. Strategies relying on training with medical datasets are highly cost-intensive and may suffer from outdated training data. Leveraging external knowledge bases is a suitable alternative, yet it faces obstacles such as limited retrieval precision and poor effectiveness in answer extraction. These issues collectively prevent LLMs from demonstrating the expected level of proficiency in mastering medical expertise. To address these challenges, we introduce Med-R2, a novel LLM physician framework that adheres to the Evidence-Based Medicine (EBM) process, efficiently integrating retrieval mechanisms as well as the selection and reasoning processes of evidence, thereby enhancing the problem-solving capabilities of LLMs in healthcare scenarios and fostering a trustworthy LLM physician. Our comprehensive experiments indicate that Med-R2 achieves an improvement of 14.74% over vanilla RAG methods and even a 3.32% enhancement compared to fine-tuning strategies, without incurring additional training costs. Furthermore, we find that our LLaMA3.1-70B + Med-R2 surpasses frontier models, including GPT-4o, Claude3.5-Sonnet and DeepSeek-V3 by 1.22%, 5.33% and 2.80%. Code and datasets supporting this study are included in our supplementary materials.
1 Introduction
Large Language Models (LLMs) have emerged as pivotal tools in the medical domain, redefining the contours of healthcare practice [1, 52, 43]. Their ability to process and understand vast amounts of unstructured medical data positions them at the forefront of medical research [8, 34], clinical decision-making [13, 27], and patient care [5, 45]. However, despite their remarkable capabilities, LLMs encounter several challenges when applied specifically to healthcare settings:
C1: Inefficiency in Knowledge Acquisition. Existing approaches predominantly rely on pretraining and fine-tuning LLMs with datasets from the medical domain, the process of which is inherently compute-intensive and resource-demanding, especially as model sizes increase [42, 32]. Moreover, training with outdated datasets can result in a lack of highly specialized expertise, which can lead to suboptimal clinical recommendations or misinform healthcare professionals [15, 25].
C2: Limited Precision for Medical Retrieval. Compared to domain-specific training, Retrieval-Augmented Generation (RAG) systems provide a cost-efficient solution, leveraging the external knowledge base to enhance content generation [28]. The retrieval quality is critical in RAG systems, where inaccuracy or misinformation can heavily influence the effectiveness of LLMs’ augmentation. While efforts [46, 19, 31] have been made to improve the retrieval precision, they neglect the specific and highly professional nature of medical knowledge, where tailored retrieval enhancement for distinct medical scenarios remains insufficiently explored.
@@ 1 Introduction
Figure 1: Comparison of Med-R2 with existing strategies for medical problem solving.
@@
C3: Low Effectiveness in Answer Extraction. Considering the constraints imposed by the models’ context window length, it is essential to critically appraise the retrieved medical evidence for its validity, impact, as well as applicability, and integrate the most pertinent ones with existing clinical expertise for problem-solving [6, 11]. Nonetheless, current studies fail to develop targeted answer extraction methods tailored for healthcare scenarios, where the nuanced evaluation of evidence hierarchies and intricate reasoning are required [40].
@@
To address these challenges, we introduce Med-R2, a noval medical LLM framework designed in accordance with the principles of Evidence-Based Medicine (EBM), conducting outstanding Retrieval and Reasoning aligned with distinct phases of EBM. 1) For C1, we have established a comprehensive external knowledge base to enhance models’ medical performances, offering a more cost-effective and flexible alternative to domain-specific training. 2) For C2, we improve the retrieval precision by refining the original queries according to their respective medical scenarios, while iteratively incorporating chain-of-thought sequences generated from the retrieved content. 3) For C3, we adopt a coarse-to-fine strategy for document appraising and filtering, and select the most pertinent ones supplemented with chain-of-thought demonstrations to assist medical queries addressing. In summary, our contributions are as follows:
@@
  • Challenges in Medical Scenarios. Through conducting a quantitative analysis of strategies aimed at enhancing models’ medical capabilities, we underscore the challenges prevalent in healthcare scenarios, including high computational consumption as well as poor efficiency in knowledge retrieval and extraction (Figure 1 and Section 4).
@@
  • LLM Physician Framework. We present Med-R2, a novel LLM physician framework that integrates the principles of evidence-based medicine (EBM) for clinical problem-solving with outstanding retrieval and reasoning, bolstering the performance of models within medical contexts (Section 3).
@@
  • Performance and Effectiveness. Our extensive experiments indicate that Med-R2 achieves a 14.74% improvement over the vanilla RAG methods, and even a 3.32% enhancement compared to the fine-tuning strategies without additional training expenses. Moreover, LLaMA3.1-70B + Med-R2 surpasses frontier models for medical problem-solving, achieving average improvements over GPT-4o, Claude3.5-Sonnet and DeepSeek-V3 by 1.22%, 5.33% and 2.80% (Section 4 and Section 5).
2 Related Work and Discussions
@@
Evidence-Based Medicine (EBM)  EBM refers to the application of the best available research to healthcare, which requires evidence integration with clinical expertise and patient values [40, 11, 39]. Clinical questions can be categorized into several types, including diagnosis, therapy, prognosis, etiology, prevention, cost, etc. [6]. Each category intersects with EBM principles by emphasizing the systematic collection, evaluation, and application of the best retrieved evidence to inform medical decision-making [12], detailed in Appendix A.
@@
Figure 2: An illustration of Med-R2’s process, adhering to the Evidence-Based Medicine (EBM) workflow. We first categorize the query by EBM and general question types. Queries are then reformulated according to established EBM classification templates to ensure precision and relevance. In the evidence searching and appraising stages, we employ a coarse-to-fine strategy to retrieve, filter, and re-rank the evidence documents within the knowledge base. CoT sequences are then generated from processed evidence to refine retrieval space, iterating to ensure robustness.
@@
LLMs for Medical Domain  As the application of LLMs expands, their deployment in the medical domain has become a widely discussed topic [8, 54]. Recent studies have concentrated on the direct use of real or synthetic medical data for the pretraining or fine-tuning of LLMs [43, 41]. Prominent open-source milestones include ChatDoctor [29] which integrates real-world doctor-patient communication data for training, PMC-LLaMA [47] pretrained on 4.9 million medical literature records, and MEDITRON [7], a scaling series of medical pretrained models. However, such extensive training can be computationally intensive. In contrast, Retrieval-Augmented Generation (RAG) systems offer a more efficient alternative, achieving comparable results with reduced training costs and enhancing the model’s precision in locating and leveraging knowledge.
@@
Retrieval-Augmented Generation (RAG)  The concept of RAG [28] was introduced as a powerful framework for integrating external knowledge into natural language generation tasks, enhancing the accuracy and relevance of generated outputs across various domains [10, 55, 18]. In the medical field, RAG has been widely used to improve LLMs’ analytical performances by utilizing external medical knowledge from sources such as medical papers, textbooks, guidelines, and entries [23, 53, 49, 50]. However, while there have been efforts dedicated to optimizing the individual components of RAG pipelines [46, 19, 4, 20], research that integrates the unique characteristics and requirements of the medical domain remains in its infancy. In this study, we incorporate the principles of Evidence-Based Medicine (EBM) into medical RAG systems to better address the special demands of healthcare.
3 Med-R2
@@
3.1 Question Formulation
Figure 3: Query category of MedMCQA. We employ a logarithmic scale (base 10) on the z-axis, ranging from 1 to 40000, to represent the wide range of values.
@@ 3.1 Question Formulation
The clinical question formulation stage consists of two components: query classifier and query reformulator. The query classification encompasses two dimensions: Evidence-Based Medicine (EBM) categories and general natural language question types. Specifically, we have delineated six distinct EBM categories and twelve general question categories, as illustrated in Figure 2. The application of this classification scheme to the MedMCQA dataset yielded the categorized results presented in Figure 3. For the classification task, we have manually annotated 100 samples to fine-tune Qwen2.5-72B-Instruct [^1] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct, and employed the trained model as our classifier. We perform domain-specific reformulations of the original queries based on their respective classes of the EBM categories to align with the professional context. Meanwhile, the general question categories are utilized as one of the criteria to rerank retrieved documents, thereby prioritizing those that best match the current query’s intent and document type preferences described in Figure 8. Further details regarding the method and implementation can be found in Section B.1.
3.2 Evidence Searching && Appraising
@@
3.2.1 Evidence Retriever
Knowledge Corpus  To construct a more comprehensive medical knowledge base adaptable to diverse healthcare scenarios, we have amassed a collection of medical data to build our retrieval corpus. The final medical knowledge base employed for retrieval comprises four distinct types of resources: academic papers, entries, books, and guidelines, with details of the data sources and statistics depicted in Section B.2.1 and Table 1. For resources with extensive content such as academic papers, books, and guidelines, we first perform content segmentation with a threshold set at 10,000 tokens. We prioritize dividing the content based on natural chapters. If natural chapters cannot be identified or exceed the threshold, we resort to truncation according to the predefined limit.
Table 1: Overall statistics of different medical knowledge resources.
@@ 3.2.1 Evidence Retriever
We integrate multiple types of retrievers to optimize our retrieval performance, incluing BGE-Large-EN-v1.5222https://huggingface.co/BAAI/bge-large-en-v1.5 for dense retrieval and SPLADE-v3333https://huggingface.co/naver/splade-v3 for sparse retrieval. We further discuss the rationale for choosing hybrid dense and sparse retrieval in Section B.2.2. We then consolidate the documents retrieved through both methods into a unified collection that contains $n$ documents, $\mathcal{D}=\left\{d_{i}\right\}_{i=1}^{n}=\mathcal{D}^{D}\cup\mathcal{D}^{S}$, where $\mathcal{D}^{D}$ comprises documents obtained by dense retrieval, and $\mathcal{D}^{S}$ represents those acquired via sparse retrieval.
3.2.2 Evidence Reranker
For the document collections retrieved in Section 3.2.1, we employ a coarse-to-fine strategy for reranking. we initially utilize the BGE-Reranker-v2-M3444https://huggingface.co/BAAI/bge-reranker-v2-m3 to rerank $\mathcal{D}=\left\{d_{i}\right\}_{i=1}^{n}$ based on their semantic relevance to the current query at a coarse granularity, returning the top $k$ documents [^5] During the previous retrieval phase, documents of all retrieval types are subjected to a unified ranking, within which the top-$k$ documents are selected for further processing., denoted as $\mathcal{S}^{c}=\left\{d_{i}\right\}_{i=1}^{k}$. At this stage, the total length of these $k$ documents significantly exceeds the context limit of the model. Then we conduct a fine-grained reranking that integrates three distinct criteria for the current $k$ documents to get $\mathcal{S}^{f}=\left\{d_{i}\right\}_{i=1}^{k}$, enabling the model to provide answers based on the most effective retrieved documents within the limited length of the context window. The fine-grained reranking score can be formulated as:
$\mathcal{F}\left(x\right)=f_{h}\left(x\right)\cdot f_{g}\left(x\right)\left(1+\alpha\cdot f_{u}\left(x\right)\right)$ (1)
@@ 3.2.2 Evidence Reranker
Hierarchy of Evidence  Medical knowledge encompasses facts and theories that are not always consistent, and sometimes even contradictory. This criterion serves dual purposes: scoring and conflict filtering. The recalled documents that have undergone coarse-grained reranking are first categorized according to their evidence levels. Subsequently, contents of these documents are analyzed for conflicting facts, where the documents that contain conflicting facts and have lower evidence ratings are filtered out. Formally, each retrieved document $d_{i}\in\mathcal{S}^{c}$ is associated with an integer evidence level $e$, where $e\in\left\{x\in\mathbb{Z}\mid 1\leq x\leq 9\right\}$, with 1 indicating the highest level of credibility:
$f_{h}\left(x\right)=9-\left(e_{x}-1\right)$ (2)
By applying the evidence assessment criteria, the retrieved documents are categorized into multiple evidence levels, and professional analysis and integration of evidence at different levels of authority are conducted, which prevents misjudgments caused by information clutter.
@@
$f_{u}\left(x\right)=\max\left\{\ell_{\theta}^{init}-\ell_{\theta}^{x},0\right\}$ (3)
@@
General Document Category  It corresponds to the mapping of general natural language question types discussed in Section 3.1. Different categories of questions desire distinct answer structures. For instance, questions regarding procedural steps are ideally answered by documents that describe processes rather than those that define concepts. To address this, we categorize the retrieved documents into 16 document types, denoted as $C$, as outlined in Table 6, and score them based on their alignment with the response type preferred by the original query. The scoring function is defined as:
$f_{g}\left(x\right)=\sum_{j=1}^{|C^{e}|}p(x|c_{j})$ (4)
@@
3.3 Evidence Applying
Through the comprehensive evaluation and reranking of retrieved documentary evidence in Section 3.2.2, we aim to ensure the application of the highest-quality available evidence to decision-making in the medical field. This process extends beyond merely applying research findings, where integrating the professional reasoning and judgment is essential. Therefore, it is imperative to make professional and reasoned inferences based on the question retrieved documents.

### v5

@@ Abstract.
Large Language Models (LLMs) have exhibited remarkable capabilities in clinical scenarios. Despite their potential, existing works face challenges when applying LLMs to medical settings. Strategies relying on training with medical datasets are highly cost-intensive and may suffer from outdated training data. Leveraging external knowledge bases is a suitable alternative, yet it faces obstacles such as limited retrieval precision and poor effectiveness in answer extraction. These issues collectively prevent LLMs from demonstrating the expected level of proficiency in mastering medical expertise. To address these challenges, we introduce Med-R2, a novel LLM physician framework that adheres to the Evidence-Based Medicine (EBM) process, efficiently integrating retrieval mechanisms as well as the selection and reasoning processes of evidence, thereby enhancing the problem-solving capabilities of LLMs in healthcare scenarios and fostering a trustworthy LLM physician. Our comprehensive experiments indicate that Med-R2 achieves an improvement of 13.27% over vanilla RAG methods and even a 4.55% enhancement compared to fine-tuning strategies, without incurring additional training costs. Furthermore, we find that our LLaMA3.1-70B + Med-R2 surpasses frontier models, including GPT-4o, Claude3.5-Sonnet and DeepSeek-V3 by 1.05%, 6.14% and 1.91%. Med-R2 effectively enhances the capabilities of LLMs in the medical domain.
Evidence-Based Medicine, Retrieval Augmented Generation, Large Language Models
copyright: acmlicensedjournalyear: 2026copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, UAE.booktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, UAE.isbn: 978-1-4503-XXXX-X/2026/06doi: XXXXXXX.XXXXXXX
@@
Figure 1. Comparison of Med-R2 with existing strategies for medical problem-solving.
1. Introduction
Large Language Models (LLMs) have emerged as pivotal tools in the medical domain, redefining the contours of healthcare practice (Abd-Alrazaq et al., 2023; Yang et al., 2023; Thirunavukarasu et al., 2023). Their ability to process and understand vast amounts of unstructured medical data positions them at the forefront of medical research (Clusmann et al., 2023; Mumtaz et al., 2024), clinical decision-making (Hager et al., 2024; Kim et al., 2024), and patient care (Busch et al., 2024; Tripathi et al., 2024). Despite their remarkable capabilities, LLMs encounter several challenges when applied specifically to healthcare settings:
C1: Inefficiency in Knowledge Acquisition. Existing approaches predominantly rely on training LLMs with datasets from the medical domain, the process of which is inherently compute-intensive and resource-demanding, especially as model sizes increase (Singhal et al., 2023b; Luo et al., 2022). Moreover, training with outdated datasets can result in a lack of highly specialized expertise, which can lead to suboptimal clinical recommendations or misinform healthcare professionals (He et al., 2022; Kandpal et al., 2023).
C2: Limited Precision for Medical Retrieval. Compared to domain-specific training, Retrieval-Augmented Generation (RAG) systems provide a cost-efficient solution, leveraging the external knowledge base to enhance content generation (Lewis et al., 2020). The retrieval quality is critical in RAG systems, where inaccuracy or misinformation can heavily influence the effectiveness of LLMs’ augmentation. While efforts (Wang et al., 2024; Jeong et al., 2024b; Long et al., 2024) have been made to improve the retrieval precision, they neglect the specific and highly professional nature of medical knowledge, where tailored retrieval enhancement for distinct medical scenarios remains insufficiently explored.
@@ 1. Introduction
C3: Low Effectiveness in Answer Extraction. Considering the constraints imposed by the models’ context window length, it is essential to critically appraise the retrieved medical evidence for its validity, impact, as well as applicability, and integrate the most pertinent ones with existing clinical expertise for problem-solving (Case and Swanson, 1998; Guyatt et al., 1992). Nonetheless, current studies fail to develop targeted answer extraction methods tailored for healthcare scenarios, where the nuanced evaluation of evidence hierarchies is required (Sackett et al., 1996).
@@
To address these challenges, we introduce Med-R2, a noval medical LLM framework designed in accordance with the principles of Evidence-Based Medicine (EBM), conducting outstanding Retrieval and Reasoning aligned with distinct phases of EBM. 1) For C1, we have established a comprehensive external knowledge base to enhance models’ medical performances, offering a more cost-effective and flexible alternative to domain-specific training. 2) For C2, we improve the retrieval precision by refining the original queries according to their respective medical scenarios, while iteratively incorporating chain-of-thought sequences generated from the retrieved content. 3) For C3, we adopt a coarse-to-fine strategy for document appraising and filtering, and select the most pertinent ones supplemented with chain-of-thought demonstrations to assist medical queries addressing. Our contributions are as follows:
@@
  • Challenges in Medical Scenarios. Through conducting a quantitative analysis of strategies aimed at enhancing models’ medical capabilities, we underscore the challenges prevalent in healthcare scenarios, including high computational consumption as well as poor efficiency in knowledge retrieval and extraction.
@@
  • LLM Physician Framework. We present Med-R2, a novel LLM physician framework that integrates the principles of evidence-based medicine (EBM) for clinical problem-solving with outstanding retrieval and reasoning capabilities within medical contexts.
@@
  • Performance and Effectiveness. Extensive experiments indicate that Med-R2 achieves a 14.74% improvement over the vanilla RAG methods, and even a 3.32% enhancement compared to the fine-tuning strategies without additional training expenses. Moreover, LLaMA3.1-70B + Med-R2 surpasses frontier models for medical problem-solving, achieving average improvements over GPT-4o, Claude3.5-Sonnet and DeepSeek-V3 by 1.22%, 5.33% and 2.80%.
2. Related Work and Discussions
@@
Evidence-Based Medicine (EBM)  EBM refers to the application of the best available research to healthcare, which requires evidence integration with clinical expertise and patient values (Sackett et al., 1996; Guyatt et al., 1992; Sackett, 1997). Clinical questions can be categorized into several types, including diagnosis, therapy, prognosis, etiology, prevention, cost, etc. (Case and Swanson, 1998). Each category intersects with EBM principles by emphasizing the collection, evaluation, and application of the best retrieved evidence to inform medical decision-making (Guyatt et al., 2000), as detailed in Appendix A.
@@
Figure 2. An illustration of Med-R2’s process, adhering to the Evidence-Based Medicine (EBM) workflow. We first categorize the query by EBM and general question types. Queries are then reformulated according to established EBM classification templates to ensure precision and relevance. In the evidence searching and appraising stages, we employ a coarse-to-fine strategy to retrieve, filter, and re-rank the evidence documents within the knowledge base. CoT sequences are then generated from processed evidence to refine retrieval space, iterating to ensure the robustness.
@@
LLMs for Medical Domain  As the application of LLMs expands, their deployment in the medical domain has become a widely discussed topic (Clusmann et al., 2023; Zeng et al., 2020). Recent studies have concentrated on the direct use of medical data for the pretraining or fine-tuning of LLMs (Thirunavukarasu et al., 2023; Singhal et al., 2023a). Prominent open-source milestones include ChatDoctor (Li et al., 2023) which integrates real-world doctor-patient communication data for training, PMC-LLaMA (Wu et al., 2024) pretrained on 4.9 million medical literature records, and MEDITRON (Chen et al., 2023), a scaling series of medical pretrained models. However, such extensive training can be computationally intensive. In contrast, Retrieval-Augmented Generation (RAG) systems offer a more efficient alternative, achieving comparable results with reduced training costs and enhancing the model’s precision in locating and leveraging knowledge.
@@
Retrieval-Augmented Generation (RAG)  The concept of RAG (Lewis et al., 2020) was introduced as a powerful framework for integrating external knowledge into natural language generation tasks, enhancing the accuracy and relevance of generated outputs across various domains (Gao et al., 2023; Zhao et al., 2024; Izacard et al., 2023). In the medical field, RAG has been widely used to improve LLMs’ analytical performances by utilizing external medical knowledge from sources such as medical papers, textbooks, guidelines, and entries (Jin et al., 2023; Zakka et al., 2024; Xiong et al., 2024a, b). However, while there have been efforts dedicated to optimizing the individual components of RAG pipelines (Wang et al., 2024; Jeong et al., 2024b; Asai et al., 2023; Jeong et al., 2024a), research that integrates the unique characteristics and requirements of the medical domain remains in its infancy. In this study, we incorporate the principles of Evidence-Based Medicine (EBM) into medical RAG systems to better address the special demands of healthcare.
3. Med-R2
@@
3.1. Question Formulation
Figure 3. Query category of MedMCQA. We employ a logarithmic scale (base 10) on the z-axis, ranging from 1 to 40000, to represent the wide range of values.
@@ 3.1. Question Formulation
The clinical question formulation stage consists of two components: query classifier and query reformulator. The query classification encompasses two dimensions: Evidence-Based Medicine (EBM) categories and general natural language question types. Specifically, we have delineated six distinct EBM categories and twelve general question categories, as illustrated in Figure 2. The application of this classification scheme to the MedMCQA dataset yielded the categorized results presented in Figure 3. For the classification task, we have manually annotated 100 samples to fine-tune Qwen2.5-72B-Instruct [^1] https://huggingface.co/Qwen/Qwen2.5-72B-Instruct, and employed the trained model as our classifier. We perform domain-specific reformulations of the original queries based on their respective classes of the EBM categories to align with the professional context. Meanwhile, the general question categories are utilized as one of the criteria to rerank retrieved documents, thereby prioritizing those that best match the current query’s intent and document type preferences described in Figure 5. Further details can be found in Section B.1.
3.2. Evidence Searching && Appraising
@@
3.2.1. Evidence Retriever
To construct a more comprehensive medical knowledge base adaptable to diverse healthcare scenarios, we have amassed a collection of medical data to build our knowledge corpus. The final medical knowledge base employed for retrieval comprises four distinct types of resources: academic papers, entries, books, and guidelines, with details of the data sources and statistics depicted in Table 1. For resources with extensive content such as academic papers, books, and guidelines, we first perform content segmentation with a threshold set at 10,000 tokens. We prioritize dividing the content based on natural chapters. If natural chapters cannot be identified or exceed the threshold, we resort to truncation according to the predefined limit.
Table 1. Overall statistics of medical knowledge resources.
@@ 3.2.1. Evidence Retriever
We integrate multiple types of retrievers to optimize our retrieval performance, incluing BGE-Large-EN-v1.5222https://huggingface.co/BAAI/bge-large-en-v1.5 for dense retrieval and SPLADE-v3333https://huggingface.co/naver/splade-v3 for sparse retrieval. We then consolidate the documents retrieved through both methods into a unified collection that contains $n$ documents, $\mathcal{D}=\left\{d_{i}\right\}_{i=1}^{n}=\mathcal{D}^{D}\cup\mathcal{D}^{S}$, where $\mathcal{D}^{D}$ comprises documents obtained by dense retrieval, and $\mathcal{D}^{S}$ represents those acquired via sparse retrieval.
3.2.2. Evidence Reranker
For the document collections retrieved in Section 3.2.1, we employ a coarse-to-fine strategy for reranking. we initially utilize the BGE-Reranker-v2-M3444https://huggingface.co/BAAI/bge-reranker-v2-m3 to rerank $\mathcal{D}=\left\{d_{i}\right\}_{i=1}^{n}$ based on their semantic relevance to the current query at a coarse granularity, returning the top $k$ documents [^5] During the previous retrieval phase, documents of all retrieval types are subjected to a unified ranking, within which the top-$k$ documents are selected for further processing., denoted as $\mathcal{S}^{c}=\left\{d_{i}\right\}_{i=1}^{k}$. At this stage, the total length of these $k$ documents significantly exceeds the context limit of the model. Then we conduct a fine-grained reranking that integrates three distinct criteria for the current $k$ documents to get $\mathcal{S}^{f}=\left\{d_{i}\right\}_{i=1}^{k^{\prime}}$, enabling the model to provide answers based on the most effective retrieved documents within the limited length of the context window. The fine-grained reranking score can be formulated as:
(1) $\mathcal{F}\left(x\right)=f_{h}\left(x\right)\cdot f_{g}\left(x\right)\left(1+\alpha\cdot f_{u}\left(x\right)\right)$
@@ 3.2.2. Evidence Reranker
Hierarchy of Evidence  Medical knowledge encompasses facts and theories that are not always consistent, and sometimes even contradictory. This criterion serves dual purposes: scoring and conflict filtering. The recalled documents that have undergone coarse-grained reranking are first categorized according to their evidence levels. Subsequently, contents of these documents are analyzed for conflicting facts, where the documents that contain conflicting facts and have lower evidence ratings are filtered out. Formally, each retrieved document $d_{i}\in\mathcal{S}^{c}$ is associated with an integer evidence level $e$, where $e\in\left\{x\in\mathbb{Z}\mid 1\leq x\leq 9\right\}$, with 1 indicating the highest level of credibility, as depicted in Figure 4:
(2) $f_{h}\left(x\right)=9-\left(e_{x}-1\right)$
By applying the evidence assessment criteria, the retrieved documents are categorized into multiple evidence levels, and professional analysis of evidence at different levels of authority is performed, preventing misjudgments caused by information clutter.
Figure 4. Hierarchy of evidence. The base of the pyramid represents the lowest quality, while the apex denotes the highest.
Figure 5. Query Document Projection.
@@
(3) $f_{u}\left(x\right)=\max\left\{\ell_{\theta}^{init}-\ell_{\theta}^{x},0\right\}$
@@
General Document Category  It corresponds to the mapping of general natural language question types discussed in Section 3.1. Different categories of questions desire distinct answer structures. For instance, questions regarding procedural steps are ideally answered by documents that describe processes rather than those that define concepts. To address this, we categorize the retrieved documents into 16 document types, denoted as $C$, as outlined in Table 6, and score them based on their alignment with the response type preferred by the original query, as depicted in Figure 5. The scoring function is defined as:
(4) $f_{g}\left(x\right)=\sum_{j=1}^{|C^{e}|}p(x|c_{j})$
@@
Output: Scores based on document category $f_{g}(S^{c})$
Define $q^{c}$: category of the query used for retrieving
Define $\vec{p}$: category probability distribution of document $d$
1:  for each document $d_{i}$ in $\mathcal{S}^{c}$ do
2:   /* Step 1: Expected Category Mapping */
4:   /* Step 2: Document Category Probability Inference */
5:   Provide category probability distribution of $d_{i}$ referring to $M_{P}$:
$\vec{p_{i}}=\{p(d_{i}|c_{j})\}_{j=1}^{|C|}\leftarrow M_{P}(d_{i})$
6:   /* Step 3: Statistics Aggregation */
8:  end for
9:  Return $f_{g}(S^{c})=\left\{f_{g}(d_{i})\right\}_{i=1}^{|S^{c}|}$
3.3. Evidence Applying
Through the comprehensive evaluation and reranking of retrieved documentary evidence in Section 3.2.2, we aim to ensure the application of the highest-quality available evidence to decision-making in the medical field. This process extends beyond merely applying research findings, where integrating the professional reasoning and judgment is essential. Therefore, it is imperative to make professional and reasoned inferences based on the retrieved documents.